# Existence of Friedrich-Wintgen Bound States in the Continuum: Cavity with a Thin Waveguide Opening

### v1

@@ 5.2 Continuity in $\delta$ and Lipschitz continuity in $\mu$ for $f_{0}$ and $f_{1}$
By neglecting the diagonal matrix $\mathrm{Diag}(\{\lambda_{m}-\mu\}_{m=0}^{M-1})$, we establish that each entry of the finite-dimensional system (56)–(57) is continuous in $\delta$ and Lipschitz continuous in $\mu$, with a Lipschitz constant of order $\mathcal{O}(h)$. For given indices $0{\leq}m_{1},m_{2}\leq{M-1}$, the entry in the $m_{1}$-th row and $m_{2}$-th column of the matrix $\mathbf{A}$ satisfies
$\Bigg{|}\sum_{j=1}^{\infty}{\bf i}\alpha_{j}(\mu)(\phi_{j},\psi_{m_{1}}(\cdot,\delta_{1}))_{\Gamma_{h}}(\psi_{m_{2}}(\cdot,\delta_{1}),\phi_{j})_{\Gamma_{h}}$
$\qquad\qquad-\sum_{j=1}^{\infty}{\bf i}\alpha_{j}(\mu)(\phi_{j},\psi_{m_{1}}(\cdot,\delta_{2}))_{\Gamma_{h}}(\psi_{m_{2}}(\cdot,\delta_{2}),\phi_{j})_{\Gamma_{h}}\Bigg{|}$
$\leq$ $n_{1}^{2}\|\frac{1}{n(\cdot,\delta_{1})}\nabla(\psi_{m_{1}}(\cdot,\delta_{1})-\psi_{m_{1}}(\cdot,\delta_{2}))\|_{[L^{2}(R_{h})]^{2}}\|\frac{1}{n(\cdot,\delta_{1})}\nabla(\psi_{m_{2}}(\cdot,\delta_{1}))\|_{[L^{2}(R_{h})]^{2}}$
$+n_{1}^{2}\|\frac{1}{n(\cdot,\delta_{1})}\nabla(\psi_{m_{2}}(\cdot,\delta_{1})-\psi_{m_{2}}(\cdot,\delta_{2}))\|_{[L^{2}(R_{h})]^{2}}\|\frac{1}{n(\cdot,\delta_{2})}\nabla(\psi_{m_{1}}(\cdot,\delta_{2}))\|_{[L^{2}(R_{h})]^{2}}.$
Hence Lemma 3.1 gives its continuity in $\delta$. Moreover, it can be verified that
(109) $\left|\sum_{j=1}^{\infty}{\bf i}(\alpha_{j}(\mu_{1})-\alpha_{j}(\mu_{2}))(\phi_{j},\psi_{m}(\cdot,\delta))_{\Gamma_{h}}(\psi_{m}(\cdot,\delta),\phi_{j})_{\Gamma_{h}}\right|\leq\mathcal{O}(h)|\mu_{1}-\mu_{2}|,$
where the constant hided in $\mathcal{O}$ is irrelevant to $\delta$. Let $\mathbf{z}_{m_{1}}(\delta,\mu)$ denote the vector
(110) $\mathbf{z}_{m_{1}}(\delta,\mu):=\sum_{j=1}^{\infty}{\bf i}\alpha_{j}(\mu)(\phi_{j},\psi_{m_{1}}(\cdot,\delta))_{\Gamma_{h}}\mathbf{w}_{j}(\delta),$
such that $\mathbf{z}_{m_{1}}(\delta,\mu)/\sqrt{1+\lambda_{m_{1}}(\delta)}$ is a restriction of the $m_{1}$-th column of the operator $\widetilde{\mathbf{H}}(\delta,\mu)$ defined as (99). Hence, $\mathbf{z}_{m_{1}}(\delta,\mu)$ is continuous in $\delta$ and Lipschitz continuous in $\mu$. Furthermore, using Lemma 5.1 and the general Sobolev inequalities, the following estimates hold:
@@
(112) $|\mathbf{z}_{m_{1}}(\delta,\mu_{1})-\mathbf{z}_{m_{1}}(\delta,\mu_{2})|$ $=\mathcal{O}(h)|\mu_{1}-\mu_{2}|.$
By similar techniques used in Lemma 5.9, we can show that the vector $\mathbf{w}_{0}(\delta)$ is also continuous in $\delta$. Finally, observe that the $(m_{1},m_{2})$-entry of $\mathbf{V}^{*}(\mathbf{I}-\mathbf{B})^{-1}\mathbf{V}$ admits the representation $\mathbf{z}_{m_{1}}^{*}\mathbf{T}(\mathbf{I}-\mathbf{T}\mathbf{H}\mathbf{T})^{-1}\mathbf{T}\mathbf{z}_{m_{2}}$, and the $m_{2}$-th column entry of $\mathbf{v}^{*}_{0,M-1,0}\mathbf{D}(\mathbf{I}-\mathbf{B})^{-1}\mathbf{V}$ can be written as $\mathbf{w}_{0}^{*}\mathbf{T}(\mathbf{I}-\mathbf{T}\mathbf{H}\mathbf{T})^{-1}\mathbf{T}\mathbf{z}_{m_{2}}$. The desired results follow from Lemmas 5.5, 5.7, and 5.9.
@@ 5.3 Existence of Friedrich-Wintgen BICs
Under Conditions A.1–A.3 and the assumption $h\ll 1$, the system (3)–(4) admits at least one BIC characterized by $\delta\approx 0$ and $\mu\approx\lambda_{M-1}(0)$.
@@
We restrict $\delta$ and $\mu$ to the intervals $[\delta_{l},\delta_{r}]$ and $[\mu_{l},\mu_{r}]$ respectively, where
(113) $\delta_{l}<0<\delta_{r},\ \mu_{l}=\lambda_{M-3}(0)+2\epsilon,\ \text{and}\ \mu_{r}=\lambda_{M}(0)-2\epsilon.$
Assume $f_{0}(\delta,\mu,h)$ and $f_{1}(\delta,\mu,h)$ are continuous in $\delta$ and $\mu$ over these intervals, and (31), (32), (37), (79), (107) and (108) are satisfied. For $h\ll 1$ with $|f_{0}(\cdot,\cdot,h)|<\epsilon$, the following inequalities hold for each fixed $\delta$:
@@
By the Intermediate Value Theorem, there exists a unique $\mu_{0}(\delta)\in(\mu_{l},\mu_{r})$ satisfying
(116) $\lambda_{M-2}(\delta)-\mu_{0}(\delta)+f_{0}(\delta,\mu_{0}(\delta),h)=0,$
where the uniqueness follows from the inequality (107). Analyzing the difference between $\mu_{0}(\delta_{1})$ and $\mu_{0}(\delta_{2})$ for $\delta_{1},\delta_{2}\in[\delta_{l},\delta_{r}]$, we obtain
$|\mu_{0}(\delta_{1})-\mu_{0}(\delta_{2})|=$ $|\lambda_{M-2}(\delta_{1})-\lambda_{M-2}(\delta_{2})+f_{0}(\delta_{1},\mu_{0}(\delta_{1}),h)-f_{0}(\delta_{2},\mu_{0}(\delta_{2}),h)|$
$\leq$ $|\lambda_{M-2}(\delta_{1})-\lambda_{M-2}(\delta_{2})|+|f_{0}(\delta_{1},\mu_{0}(\delta_{2}),h)-f_{0}(\delta_{2},\mu_{0}(\delta_{2}),h)|$
$+|f_{0}(\delta_{1},\mu_{0}(\delta_{1}),h)-f_{0}(\delta_{1},\mu_{0}(\delta_{2}),h)|$
$\leq$ $|\lambda_{M-2}(\delta_{1})-\lambda_{M-2}(\delta_{2})|+|f_{0}(\delta_{1},\mu_{0}(\delta_{2}),h)-f_{0}(\delta_{2},\mu_{0}(\delta_{2}),h)|$
(117) $+C_{f_{0}}|\mu_{0}(\delta_{1})-\mu_{0}(\delta_{2})|.$
This implies the continuity of $\mu_{0}(\delta)$ since $C_{f_{0}}<1$. Similarly, there exists a unique continuous function $\mu_{1}(\delta)$ satisfying
(118) $\lambda_{M-1}(\delta)-\mu_{1}(\delta)+f_{1}(\delta,\mu_{1}(\delta),h)=0.$
Without loss of generality, we assume $\lambda_{M-2}(\delta_{r})>\lambda_{M-1}(\delta_{r})$, $\lambda_{M-2}(\delta_{l})<\lambda_{M-1}(\delta_{l})$ and $h\ll 1$ such that
(119) $|f_{0}(\cdot,\cdot,h)|+|f_{1}(\cdot,\cdot,h)|\leq\min\Big{\{}\lambda_{M-2}(\delta_{r})-\lambda_{M-1}(\delta_{r}),\lambda_{M-1}(\delta_{l})-\lambda_{M-2}(\delta_{l})\Big{\}}.$
From this bound, we deduce
(120) $\mu_{0}(\delta_{r})-\mu_{1}(\delta_{r})\geq 0,\ \mu_{0}(\delta_{l})-\mu_{1}(\delta_{l})\leq 0.$
By the Intermediate Value Theorem, $\mu_{0}(\delta)$ and $\mu_{1}(\delta)$ must coincide for at least one $\delta$.
@@ 6 Numerical experiments
The refractive index $n(\mathbf{x})$ is modeled as a piecewise constant function: it assumes the same constant value within each element and is set to $1$ in the surrounding ambient medium.
@@
For the eigenvalue problem (5)–(6) in the region $\Omega_{\mathrm{in}}$, Figure 3 illustrates the evolution of the first eight eigenvalues as the refractive index $n$ of the elements increases from $1$ to $2$. Notably, the sixth and seventh eigenvalues, labeled $\lambda_{6}$ and $\lambda_{7}$, cross at $n\approx 1.461$, resulting in a multiple eigenvalue $\lambda_{6}=\lambda_{7}\approx 1.695$.
When a thin waveguide is introduced on the boundary of $\Omega_{\mathrm{in}}$, the resulting resonances approximating $\lambda_{6}$ and $\lambda_{7}$ are denoted by $\widetilde{\lambda}_{6}$ and $\widetilde{\lambda}_{7}$. Their real and imaginary parts are plotted in Fig. 4(a) and 4(b), respectively. A BIC emerges at $n\approx 1.442$ and $\widetilde{\lambda}_{6}\approx 1.718$, confirming our theoretical predictions.
@@
Figure 4: Resonances of the problem (3)–(4) in $\Omega$ approximating the eigenvalues $\lambda_{6}$ and $\lambda_{7}$ as the refractive index of the elements varies.
@@ 7 Discussion and conclusion
The scaling transformation considered in [16] is equivalent to defining $\bm{\mathcal{R}}$ as:
@@
where $\lambda$ is a simple eigenvalue associated with an odd eigenfunction and $a(\mu,h)$ represents a function depending on $\mu$ and $h$. By demonstrating the uniform convergence of $a(\cdot,h)$ to 00 as $h\to 0$ and the continuity of $a(\mu,h)$ in $\mu$, the existence of a BIC is guaranteed for $h\ll 1$ by the Intermediate Value Theorem.

### v2

@@ 5.2 Continuity in $\delta$ and Lipschitz continuity in $\mu$ for $f_{0}$ and $f_{1}$
By neglecting the diagonal matrix $\mathrm{Diag}(\{\lambda_{m}-\mu\}_{m=0}^{M-1})$, we will show that each entry of the finite-dimensional system (58)–(59) is continuous in $\delta$ and Lipschitz continuous in $\mu$, with a Lipschitz constant of order $\mathcal{O}(h)$. The desired result follows from this. Consider the $(m_{1},m_{2})$-entry of $\mathbf{A}$ for $0{\leq}m_{1},m_{2}\leq{M-1}$. Its difference for two values of $\delta$ is bounded by
$\Bigg|\sum_{j=1}^{\infty}{\bf i}\alpha_{j}(\mu)(\phi_{j},\psi_{m_{1}}(\cdot,\delta_{a}))_{\Gamma_{h}}(\psi_{m_{2}}(\cdot,\delta_{a}),\phi_{j})_{\Gamma_{h}}$
$\qquad\qquad-\sum_{j=1}^{\infty}{\bf i}\alpha_{j}(\mu)(\phi_{j},\psi_{m_{1}}(\cdot,\delta_{b}))_{\Gamma_{h}}(\psi_{m_{2}}(\cdot,\delta_{b}),\phi_{j})_{\Gamma_{h}}\Bigg|$
$\leq$ $n_{1}^{2}\|\frac{1}{n(\cdot,\delta_{a})}\nabla(\psi_{m_{1}}(\cdot,\delta_{a})-\psi_{m_{1}}(\cdot,\delta_{b}))\|_{[L^{2}(R_{h})]^{2}}\|\frac{1}{n(\cdot,\delta_{a})}\nabla(\psi_{m_{2}}(\cdot,\delta_{a}))\|_{[L^{2}(R_{h})]^{2}}$
$+n_{1}^{2}\|\frac{1}{n(\cdot,\delta_{a})}\nabla(\psi_{m_{2}}(\cdot,\delta_{a})-\psi_{m_{2}}(\cdot,\delta_{b}))\|_{[L^{2}(R_{h})]^{2}}\|\frac{1}{n(\cdot,\delta_{b})}\nabla(\psi_{m_{1}}(\cdot,\delta_{b}))\|_{[L^{2}(R_{h})]^{2}}.$
By Lemma 3.1, this entry is continuity in $\delta$. And its Lipschitz continuity in $\mu$ follows from Lemmas 4.3, 5.1 and the general Sobolev inequality, which give
(113) $\left|\sum_{j=1}^{\infty}{\bf i}(\alpha_{j}(\mu_{a})-\alpha_{j}(\mu_{b}))(\phi_{j},\psi_{m}(\cdot,\delta))_{\Gamma_{h}}(\psi_{m}(\cdot,\delta),\phi_{j})_{\Gamma_{h}}\right|\leq\mathcal{O}(h^{2})|\mu_{a}-\mu_{b}|,$
where the constant hided in $\mathcal{O}$ is independent of $\delta$. Let the vector $\mathbf{z}_{m_{1}}(\delta,\mu)$ be defined by
(114) $\mathbf{z}_{m_{1}}(\delta,\mu):=\sum_{j=1}^{\infty}{\bf i}\alpha_{j}(\mu)(\phi_{j},\psi_{m_{1}}(\cdot,\delta))_{\Gamma_{h}}\mathbf{w}_{j}(\delta).$
The scaled vector $\mathbf{z}_{m_{1}}(\delta,\mu)/\sqrt{1+\lambda_{m_{1}}(\delta)}$ is a restriction of the $m_{1}$-th column of the operator $\widetilde{\mathbf{H}}(\delta,\mu)$ from (102). Applying Lemma 5.1 and the general Sobolev inequality yields the estimates:
@@
(116) $\|\mathbf{z}_{m_{1}}(\delta,\mu_{1})-\mathbf{z}_{m_{1}}(\delta,\mu_{2})\|_{\ell^{2}}$ $=\mathcal{O}(h)|\mu_{1}-\mu_{2}|.$
Using techniques analogous to those in Lemma 5.9, we also find that the vector $\mathbf{w}_{0}(\delta)$ is continuous in $\delta$. The proof is completed by observing that the $(m_{1},m_{2})$-entry of $\mathbf{V}^{*}(\mathbf{I}-\mathbf{B})^{-1}\mathbf{V}$ admits the representation $\mathbf{z}_{m_{1}}^{*}\mathbf{T}(\mathbf{I}-\mathbf{T}\mathbf{H}\mathbf{T})^{-1}\mathbf{T}\mathbf{z}_{m_{2}}$, and the $m_{2}$-th column entry of $\mathbf{v}^{*}_{0,M-1,0}\mathbf{D}(\mathbf{I}-\mathbf{B})^{-1}\mathbf{V}$ can be written as $\mathbf{w}_{0}^{*}\mathbf{T}(\mathbf{I}-\mathbf{T}\mathbf{H}\mathbf{T})^{-1}\mathbf{T}\mathbf{z}_{m_{2}}$. Their continuity in $\delta$ and Lipschitz continuity in $\mu$ follow from Lemmas 5.5, 5.7, and 5.9. And the value $h_{3}$ is chosen such that (111)–(112) hold for all $h\in(0,h_{3})$.
@@ 5.3 Existence of Friedrich-Wintgen BICs
Under Conditions A.1–A.3, there exists $h_{4}>0$ such that for every $h\in(0,h_{4})$, the system (3)–(4) admits a BIC for some $\delta\in[-\delta_{3},\delta_{3}]$ and $\mu\in[\mu_{l},\mu_{r}]$, where $\delta$ and $\mu$ depend on $h$.
@@
Let $\delta_{l}:=-\delta_{3}$ and $\delta_{r}:=\delta_{3}$. By Theorem 5.3, for sufficiently small $h$, we have
(117) $|f_{0}(\delta,\mu,h)|<\epsilon\ \text{and}\ |f_{1}(\delta,\mu,h)|<\epsilon\ \text{for}\ \delta\in[\delta_{l},\delta_{r}],\ \mu\in[\mu_{l},\mu_{r}].$
For a fixed $\delta$, this implies the inequalities:
@@
Since $f_{0}$ is continuous in $\mu$ by Theorem 5.11, the Intermediate Value Theorem guarantees the existence of $\mu_{0}(\delta)\in(\mu_{l},\mu_{r})$ satisfying
(120) $\lambda_{M-2}(\delta)-\mu_{0}(\delta)+f_{0}(\delta,\mu_{0}(\delta),h)=0.$
And the Lipschitz condition (111) ensures $\mu_{0}$ is uniquely defined. The continuity of $\mu_{0}(\delta)$ follows from estimating the difference for $\delta_{a},\delta_{b}\in[\delta_{l},\delta_{r}]$:
$|\mu_{0}(\delta_{a})-\mu_{0}(\delta_{b})|=$ $|\lambda_{M-2}(\delta_{a})-\lambda_{M-2}(\delta_{b})+f_{0}(\delta_{a},\mu_{0}(\delta_{a}),h)-f_{0}(\delta_{b},\mu_{0}(\delta_{b}),h)|$
$\leq$ $|\lambda_{M-2}(\delta_{a})-\lambda_{M-2}(\delta_{b})|+|f_{0}(\delta_{a},\mu_{0}(\delta_{b}),h)-f_{0}(\delta_{b},\mu_{0}(\delta_{b}),h)|$
$+|f_{0}(\delta_{a},\mu_{0}(\delta_{a}),h)-f_{0}(\delta_{a},\mu_{0}(\delta_{b}),h)|$
$\leq$ $|\lambda_{M-2}(\delta_{a})-\lambda_{M-2}(\delta_{b})|+|f_{0}(\delta_{a},\mu_{0}(\delta_{b}),h)-f_{0}(\delta_{b},\mu_{0}(\delta_{b}),h)|$
(121) $+C_{f_{0}}|\mu_{0}(\delta_{a})-\mu_{0}(\delta_{b})|.$
Since $C_{f_{0}}<1$, this inequality implies the continuity of $\mu_{0}(\delta)$. A similar argument shows the existence of a unique, continuous function $\mu_{a}(\delta)$ satisfying
(122) $\lambda_{M-1}(\delta)-\mu_{a}(\delta)+f_{1}(\delta,\mu_{a}(\delta),h)=0.$
Invoking Theorem 5.3, we further take $h$ sufficiently small to satisfy
$|f_{0}(\cdot,\cdot,h)|+|f_{1}(\cdot,\cdot,h)|$
(123) $\leq$ $\min\Big\{|\max_{\delta\in[\delta_{l},\delta_{r}]}\left\{\lambda_{M-2}(\delta)-\lambda_{M-1}(\delta)\right\}|,\ |\min_{\delta\in[\delta_{l},\delta_{r}]}\left\{\lambda_{M-2}(\delta)-\lambda_{M-1}(\delta)\right\}|\Big\}.$
From this and Condition A.3, it follows that
(124) $\max_{\delta\in[\delta_{l},\delta_{r}]}\{\mu_{0}(\delta)-\mu_{1}(\delta)\}\geq 0\ \text{and}\ \min_{\delta\in[\delta_{l},\delta_{r}]}\{\mu_{0}(\delta)-\mu_{1}(\delta)\}\leq 0.$
Therefore, the continuous function $\mu_{0}(\delta)-\mu_{1}(\delta)$ must have a zero in $\delta_{l},\delta_{r}]$ by the Intermediate Value Theorem. We may thus choose $h_{4}>0$ such that (117) and (123) are satisfied for all $h\in(0,h_{4})$.
@@ 6.1 Example 1
The refractive index $n(\mathbf{x})$ is modeled as a piecewise constant function: it assumes the same constant value within each element and is set to $1$ in the surrounding ambient medium. We use a Perfectly Matched Layer (PML) to absorb outgoing fields and employ the piecewise linear continuous finite-element method to formulate the eigenvalue problems. Fig. 3 and 3 show examples of our computational meshes, all of which are constructed uniformly.
@@
Figure 3: (a). Computational mesh used for the eigenvalue problem (5)–(6) in $\Omega_{\mathrm{in}}$. (b) Computational mesh used for computing the resonances of the system governed by (3)–(4) in $\Omega$ with $h=2\pi/9$.
For the eigenvalue problem (5)–(6) in the region $\Omega_{\mathrm{in}}$, Figure 4 illustrates the evolution of the first eight eigenvalues as the refractive index $n$ of the elements increases from $1$ to $2$. Notably, the sixth and seventh eigenvalues, labeled $\lambda_{6}$ and $\lambda_{7}$, intersect transversally at $n\approx 1.461$, resulting in a multiple eigenvalue $\lambda_{6}=\lambda_{7}\approx 1.695$.
@@
When a thin waveguide is introduced on the boundary of $\Omega_{\mathrm{in}}$, the resulting resonances approximating $\lambda_{6}$ and $\lambda_{7}$ are denoted by $\widetilde{\lambda}_{6}$ and $\widetilde{\lambda}_{7}$. Their real and imaginary parts are plotted in Fig. 5(a) and 5(b), respectively. A BIC emerges at $n\approx 1.442$ and $\widetilde{\lambda}_{6}\approx 1.718$, confirming our theoretical predictions. The corresponding BIC mode is shown in Fig. 5(c).
@@
(c) The BIC mode at $n\approx 1.442$ and $\mu=\widetilde{\lambda}_{6}\approx 1.718$.
Figure 5: Resonances of the problem (3)–(4) in $\Omega$ approximating the eigenvalues $\lambda_{6}$ and $\lambda_{7}$ as the refractive index of the elements varies. A BIC emerges at $n\approx 1.442$ and $\mu=\widetilde{\lambda}_{6}\approx 1.718$.
6.2 Example 2
We further demonstrate the robustness of the FW-BICs with an example where the waveguide width is increased to $h=4\pi/9$. The computational mesh for the finite-element simulation is shown in Fig. 6. The resulting resonances approximating $\lambda_{6}$ and $\lambda_{7}$ are denoted by $\widehat{\lambda}_{6}$ and $\widehat{\lambda}_{7}$, whose real and imaginary parts are plotted in Fig. 7(a) and 7(b), respectively. A BIC emerges at $n\approx 1.385$ and $\widetilde{\lambda}_{6}\approx 1.771$, and its corresponding mode profile is shown in Fig. 7(c).
Figure 6: Computational mesh used for computing the resonances of the system governed by (3)–(4) in $\Omega$ with $h=4\pi/9$.
(a) Real parts of $\widehat{\lambda}_{6}$ and $\widehat{\lambda}_{7}$.
(b) Imaginary parts of $\widehat{\lambda}_{6}$ and $\widehat{\lambda}_{7}$.
(c) The BIC mode at $n\approx 1.385$ and $\mu=\widehat{\lambda}_{6}\approx 1.771$.
Figure 7: Resonances of the problem (3)–(4) in $\Omega$ approximating the eigenvalues $\lambda_{6}$ and $\lambda_{7}$ as the refractive index of the elements varies. A BIC emerges at $n\approx 1.385$ and $\mu=\widehat{\lambda}_{6}\approx 1.771$.
@@ 7 Discussion and conclusion
[…]
@@
where $\lambda$ is a simple eigenvalue associated with an odd eigenfunction and $a(\mu,h)$ represents a function depending on $\mu$ and $h$. By demonstrating the uniform convergence of $a(\cdot,h)$ to 0 as $h\to 0$ and the continuity of $a(\mu,h)$ in $\mu$, the existence of a BIC is guaranteed for $h\ll 1$ by the Intermediate Value Theorem.